OpenStudy (anonymous):

can someone help me out please? solve the system by substitution. y=x-1 y=-2x+5

jigglypuff314 (jigglypuff314):

we know that y equals y right? therefore you can get that x - 1 = -2x + 5 then solve for x, what do you get? :)

OpenStudy (anonymous):

umm y= (-2x+5) -1? would I start by doing that?

jigglypuff314 (jigglypuff314):

mmm you would start by solving for x from x - 1 = -2x + 5

OpenStudy (anonymous):

ok... wouldn't that come out as 2?

jigglypuff314 (jigglypuff314):

yep! :D so now then to get the y value you plug in what you got for x back into one of the original equations. so plug in 2 for x into y = x - 1

OpenStudy (anonymous):

so y= (2)-1 =1 y=1?

jigglypuff314 (jigglypuff314):

yep! :D

OpenStudy (anonymous):

cool! ok so now would I plug 2 into y=-2x+5 or something, or am I done?

jigglypuff314 (jigglypuff314):

you can do that to double check if you would like :) but otherwise you should be done. unless your teacher wants the answer in ordered pair form which would be (x, y) -> (2, 1)

OpenStudy (anonymous):

it doesn't ask for that so sweet, looks like you helped me finish the problem! thank you!

jigglypuff314 (jigglypuff314):

glad I could help :)